THE PROBLEMS ARE SEPERATED BY |

1/3 x 1/4 | 13/16 x 1/3 | 5/6 x 7/8
1
Expert's answer
2012-09-13T08:14:20-0400
1/3 x 1/4=1/(3*4)=1/12
13/16 x 1/3=13/(16*3)=13/48
5/6 x 7/8=(5*7)/(6*8)=35/48